y'' + y' - 6y = 2x

1
Expert's answer
2021-03-12T12:18:52-0500

Solution.

"y''+y'-6y=2x."


This is a linear differential equation of the second-order with constant coefficients. The general solution of this equation is found as the sum of the general solution "\tilde{y}" ​ of the corresponding homogeneous equation and some particular integral solution "y_*\u200b" of the inhomogeneous equation:"y=\tilde{y}+y_*."

Сompose and solve the characteristic equation:


"\lambda^2+\lambda-6=0, \newline
\lambda_1=-3, \lambda_2=2."

 Therefore, "\tilde{y}=C_1e^{\lambda_1x}+C_2e^{\lambda_2x}." So,

"\tilde{y}=C_1e^{-3x}+C_2e^{2x}."

Since the right-hand side also contains a "2x" , we find a particular integral solution "y_*" ​ ​in the form: 

"y_*=Ax+B."


Using the method of undetermined coefficients we find that "A=-\frac{1}{3}, B=-\frac{1}{18}."

So, the particular integral solution of equation "y''+y'-6y=2x" is

"y_*=-\frac{1}{3}x-\frac{1}{18}."


And the general solution of the same equation is


"y=C_1e^{-3x}+C_2e^{2x}-\frac{1}{3}x-\frac{1}{18}."

Answer. "y=C_1e^{-3x}+C_2e^{2x}-\frac{1}{3}x-\frac{1}{18}."